OpenStudy (anonymous):

A hot air balloon is at a height of h feet. Which expressions represent a change in height of 5 ft? Choose exactly two answers that are correct. A. h/5 B. 5h C. h+5 D. h-5

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@acxbox22

OpenStudy (anonymous):

@gottennis121

pooja195 (pooja195):

what do you think the word change means?

OpenStudy (acxbox22):

a change in this context can only go up or down so the height can only go 2 ways which 2 ways are they?

OpenStudy (anonymous):

i think it is B and C

OpenStudy (acxbox22):

you are at your current height of h now you have a CHANGE in height the CHANGE can be either negative or positive

OpenStudy (anonymous):

negative

OpenStudy (anonymous):

so B and D

OpenStudy (anonymous):

@acxbox22

OpenStudy (acxbox22):

hint:height can only go up or go down for example you are at an altitude of 20 ft you have a change in altitude of 10 ft so you can either go up 10 ft or go down 10 ft going up = 20+10 going down=20-10

OpenStudy (gottennis121):

I guess you got help from @acxbox22, and you don't need my help @Deadpool117. :D

OpenStudy (anonymous):

B and D

OpenStudy (anonymous):

or B and C

OpenStudy (acxbox22):

D is correct but B is wrong remember you are not multiplying your change by your height you are adding

OpenStudy (anonymous):

so A is my secod answer

OpenStudy (anonymous):

@acxbox22

OpenStudy (anonymous):

he keeps leaving .-.

OpenStudy (acxbox22):

no i am here ;) dont worry

OpenStudy (acxbox22):

A would be dividing our height by our change that is not correct this only leaves us with 2 correct answer choices

OpenStudy (anonymous):

C and D

OpenStudy (acxbox22):

\(\Huge\color{green}{correct!}\)